OpenStudy (anonymous):

Find antiderivative of 6/7csc^2 x/7

OpenStudy (anonymous):

the derivative of csc^2(x) is -cot(x) . hope this helps

OpenStudy (anonymous):

skipping steps: = 6/7 * -cot(x/7) * 7 = - 6 cot(x/7) let me know if you have questions

OpenStudy (anonymous):

+ constant ^_^

OpenStudy (anonymous):

I know that csc^2x equals -cot(x) but how did you solve it

jimthompson5910 (jim_thompson5910):

Euler271 meant to say the derivative of cot(x) is -csc^2(x) or the derivative of -cot(x) is csc^2(x)

jimthompson5910 (jim_thompson5910):

the trick is to use u-sub let u = x/7 so du/dx = 1/7 6du/dx = 6/7 6du = (6/7)dx (6/7)dx = 6du

OpenStudy (anonymous):

but the 7's wouldn't cancel out

jimthompson5910 (jim_thompson5910):

what do you mean

OpenStudy (anonymous):

the seven in (6/7) and the 1/7, does not cancel out the sevens.. giving me -6cot(x/7) + C soo... how does the sevens cancel out in your answer?

jimthompson5910 (jim_thompson5910):

I didn't cancel out the 7s though

jimthompson5910 (jim_thompson5910):

do you see how I got du/dx = 1/7

OpenStudy (anonymous):

sorry but no, please explain in detail

jimthompson5910 (jim_thompson5910):

ok I let u = x/7 because this is the argument of csc^2

jimthompson5910 (jim_thompson5910):

then I derived both sides to get du/dx = 1/7

jimthompson5910 (jim_thompson5910):

after that I multiplied both sides by 6 to get 6du/dx = 6/7

jimthompson5910 (jim_thompson5910):

and then I multiplied both sides by dx to get 6du = (6/7)dx

jimthompson5910 (jim_thompson5910):

So if we have this